OpenStudy (anonymous):

a compounds percent composition is 43.6% P and 56.4% O. What is the empirical formula of the compound?

OpenStudy (anonymous):

P: 43.6 / 30.97= 1.40 1.40/1.40= 1 x 2 ( to make it a whole number) = 2 O: 56.4/16= 3.52 3.45/1.40 = 2.4 x2 ( to make it a whole number) = 8.2 Empirical Formula P2O8